# JoIN: Joint GANs Inversion for Intrinsic Image Decomposition

### v2

@@ Abstract
In this work, we propose to solve ill-posed inverse imaging problems using a bank of Generative Adversarial Networks (GAN) as a prior and apply our method to the case of Intrinsic Image Decomposition for faces and materials. Our method builds on the demonstrated success of GANs to capture complex image distributions. At the core of our approach is the idea that the latent space of a GAN is a well-suited optimization domain to solve inverse problems. Given an input image, we propose to jointly inverse the latent codes of a set of GANs and combine their outputs to reproduce the input. Contrary to most GAN inversion methods which are limited to inverting only a single GAN, we demonstrate that it is possible to maintain distribution priors while inverting several GANs jointly. We show that our approach is modular, allowing various forward imaging models, and that it can successfully decompose both synthetic and real images.
@@ 1 Introduction
Figure 1: Overview of our approach. We train multiple GANs, each trained only on a single image component, their output is combined using a function $f$. The latent codes of each GAN are optimized to minimize the loss between the target image $I^{*}$ and the output $I$.
@@
Any natural image can be seen as an intricate combination of geometry-dependent, material-dependent, and lighting-dependent components. Decomposing an image in such components, a problem known as intrinsic image decomposition (IID) allows for varying each property independently in order to achieve complex image editing [13].
While the forward mapping from the decomposed image components to the natural image is direct, the inverse process is a challenging ill-posed problem. To obtain a useful decomposition it is usual to constrain the solution space by employing either hand-crafted [69, 5] or learned priors [47]. However, these priors are often restricting the solution too much or lead to cross-contamination where texture is copied in shading and albedo exhibits shading variations.
Recently, there has been a growing interest in employing Generative Adversarial Networks (GANs) as priors in a variety of inverse imaging problems such as image inpainting [78, 12, 27], HDR imaging [56], image super-resolution [44, 12, 27], compressive sensing [12, 66], and phase retrieval [29, 33], owing to their tremendous success in mimicking real data distributions. A common way to leverage these GAN priors is to invert the mapping of the GAN – a highly challenging problem known as GAN Inversion. While existing GAN Inversion methods are limited to inverting only a single GAN, in this work we propose to jointly invert multiple GANs to decompose natural image. Contrary to previous work in which all components are known in advance and have to be generated by the same GAN [28], we propose to train multiple GANs independently. We show this prevents channel cross-contamination and also allows flexible forward and inverse models as the number of components and the way they are combined does not have to be pre-determined.
We apply this new approach to intrinsic image decomposition by training a bank of GANs for components such as albedo or shading. During inversion, to ensure that each intrinsic component remains on the distribution manifold of its associated GAN we propose a new kNN-based loss regularization to constrain the optimization. We further demonstrate that both encoder-based GAN initialization and generator tuning techniques from the existing literature can be used on this joint inversion problem, improving the results. We test and evaluate our method on two types of datasets, faces and materials. To summarize, we introduce a novel framework to solve image-based inverse problems which we apply to IID with the following key contributions:
  • We propose to use a bank of GANs as a prior to solve inverse problems by jointly inverting them.
@@
  • We introduce a novel kNN-based regularization term for GAN inversion allowing them to better retain their distribution properties than existing in-domain loss.
  • We show that this approach is both flexible and compatible with existing GAN inversion techniques.
@@ 2.1 Image Decomposition
The intrinsic decomposition of images is a long-standing problem that emerged with the retinex and lightness theory [42, 30]. Using a purely diffuse model, early work computed shadow-free reflectance images of faces [74] using time-lapse sequences. User-assisted methods were the first to successfully tackle the problem for single natural images  [70, 13]. With the rise of deep learning a large body of work emerged using neural networks to predict intrinsic images [83, 55, 46, 45, 80, 47]. Recently, Liu et al. [50] presented a method, sharing some ideas with our pipeline. They build priors independently for shading and reflectance using image collections and adversarial auto-encoders. Contrary to our approach, they use a fixed forward model and would require full retraining if a component such as specular were to be added.
@@
For photo collections, many works explored this problem using geometry priors [40, 41, 24, 59, 60], and recently methods were introduced working on videos [77, 11, 52, 53]. One of the main applications of IID lies in image relighting for which decomposing an image in specific components such as albedo, shading, shadows or normals is often an intermediate step both for faces [57, 31, 79] and outdoor picture [26]. In the context of the more constrained problem of material acquisition, methods often aim at recovering richer information. They usually predict a full SVBRDF map composed of albedo but also normals, roughness, and specular maps [3, 19, 20]. Some recent methods proposed to use adversarial networks as part of their SVBRDF estimation pipelines [84, 85]. Closer in spirit to our method MaterialGAN [28] proposes to use a single GAN trained on all the channels of SVBRDF maps and to invert a latent code to recover an SVBRDF. Contrary to our method it requires several input images and known lighting conditions for each input.
@@ 2.2 GAN Priors and GAN Inversion
Inverting a GAN to obtain the latent code corresponding to an image is a key step in solving inverse problems [67]. Owing to its importance, many GAN inversion approaches have been recently proposed as reviewed in [75]. We can classify these methods in three broad classes: optimization-based, encoder-based, and hybrid.
@@
Optimization-based methods such as [48, 17, 86, 49, 38, 1, 2] regress a latent code using gradient descent by minimizing a loss between a target image and the generated image produced by the latent code. The key variations among these approaches are the latent space chosen for the optimization ($\mathcal{Z}$ space, $\mathcal{W}$ space, $\mathcal{W^{+}}$ space, etc.); the loss function(s) they use to calculate the similarity between the target and the estimate; and additional criterion used for aiding optimization (such as regularization, improved initialization, and early stopping). Although these methods can produce relatively precise outcomes, their iterative optimization scheme is slower than forward methods and can converge to local minima when badly initialized.
@@
Contrary to optimization-based approaches, encoder-based methods propose to use an end-to-end framework based in which an encoder that takes the target image as input, predicts the latent code directly [75, 61, 63, 4, 15, 43, 86, 14, 58, 71, 73, 72, 76, 54, 51]. These different methods propose powerful encoder architectures, specialized loss functions, and training procedures adapted to GAN inversion. The main advantage of encoder-based methods is their faster inference times, but they generally suffer from lower result quality and worse generalization. This makes them great candidates to serve as an initializer for optimization-based inversion. Methods that combine both optimization and encoder-based initialization are typically known as Hybrid approach [16, 8, 86, 4, 7, 6, 32, 73].
@@
Recently, a fourth category of approaches has emerged that advocates fine-tuning the generator weights after optimizing the latent code [65, 25]. Such methods produce better reconstruction since the generator is allowed to fit the target image but fine-tuning the generator can lead to distortions in the generator manifold losing important properties. While our proposed method and contributions are orthogonal to the different inversion approaches we show that they are compatible with joint GAN inversion. More recently, [9] has shown that StyleGAN has an internal representation of intrinsic images that can be accessed by finding an appropriate offset to the latent codes. However, this method is not applicable for real images, since the offsets obtained are not compatible with the output latent codes of the GAN inversion on the real image.
@@ 3 Our Approach
In this work, we propose to solve intrinsic image decomposition using an optimization formulation that leverages a bank of GANs as priors. Our aim is to decompose a natural image into its material-dependent and lighting-dependent components such as albedo and shading. This inverse problem is ill-posed, and priors on the structure of each image component are required to obtain a solution.
Each image component distribution can be closely approximated using a GAN model. We propose to create a bank of GANs that consists of one GAN model for each image component that we desire to recover. Since each GAN maps a lower dimensional latent space to an image component space, solving the inverse problem aims to recover the latent codes corresponding to each image component which can be formulated as the joint inversion of multiple GANs.
@@ 3.1 JoIN: Joint Inversion of Multiple GANs
For $G_{i}$ we follow the architecture of StyleGAN2 [37] where $w_{i}$ is in the $\mathcal{W}$ space. Our estimate of $I$ is thus given by:
@@
Eq. 4 is analogous to the standard GAN inversion formulation, except for the fact that it aims to invert $n$ Generators simultaneously instead of one. The loss in Eq. 4 is minimized over all the latent codes $w_{i}$s simultaneously, and the final estimates $\widehat{w}_{i}$s can be passed to Eq. 2 to obtain the image components $C^{*}_{1\dots n}$. We use E-LPIPS loss [39] (robust version of LPIPS [82] perceptual loss) as the loss function $\mathcal{L}$ unless otherwise specified.

### v3

@@ Abstract
Intrinsic Image Decomposition (IID) is a challenging inverse problem that seeks to decompose a natural image into its underlying intrinsic components such as albedo and shading. While recent image decomposition methods rely on learning-based priors on these components, they often suffer from component cross-contamination owing to joint training of priors; or from Sim-to-Real gap since the priors trained on synthetic data are kept frozen during the inference on real images. In this work, we propose to solve the intrinsic image decomposition problem using a bank of Generative Adversarial Networks (GANs) as priors where each GAN is independently trained only on a single intrinsic component, providing stronger and more disentangled priors. At the core of our approach is the idea that the latent space of a GAN is a well-suited optimization domain to solve inverse problems. Given an input image, we propose to jointly invert the latent codes of a set of GANs and combine their outputs to reproduce the input. Contrary to all existing GAN inversion methods that are limited to inverting only a single GAN, our proposed approach, JoIN, is able to jointly invert multiple GANs using only a single image as supervision while still maintaining distribution priors of each intrinsic component. We show that our approach is modular, allowing various forward imaging models, and that it can successfully decompose both synthetic and real images. Further, taking inspiration from existing GAN inversion approaches, we allow for careful fine-tuning of the generator priors during the inference on real images. This way, our method is able to achieve excellent generalization on real images even though it uses only synthetic data to train the GAN priors. We demonstrate the success of our approach through exhaustive qualitative and quantitative evaluations and ablation studies on various datasets.
@@ 1 Introduction
Any natural image can be seen as an intricate combination of geometry-dependent, material-dependent, and lighting-dependent components. Decomposing an image into such intrinsic components is highly useful, since it allows for varying each intrinsic component independently in order to achieve complex image editing such as image relighting, material swap, and object insertion (Bousseau et al., 2009).
In a commonly used image decomposition framework, an image is typically decomposed into a light independent component referred to as albedo, a light dependent component referred to as shading, and optionally a residual component referred to as specular which captures the non-ideal behavior of natural surfaces. As depicted in Fig. 1, the forward mapping $f(\cdot)$ from the decomposed image components to the natural image in this case is relatively straightforward to obtain. However, the inverse mapping of the natural image to its intrinsic components is highly challenging ill-posed inverse problem owing to potentially infinite solution space. To obtain useful decomposition, it is required to constrain the solution space by employing priors on the image components.
Existing IID approaches leverage variety of hand-crafted or learned priors for recovery of different intrinsic components. However, priors from both of these categories often suffer from insufficient disentanglement and cross-contamination of components, i.e. features from shading may appear in albedo and vice versa. This effect is more pronounced in learning-based approaches that attempt to train a single model to predict multiple image components jointly (Das et al., 2022a; Yu & Smith, 2021). Further, in the case of learning-based priors, collecting the training data with groundtruth albedo and shading of natural images is extremely difficult, thus, they are typically trained using the simulated/synthetic data. Priors trained on simulated data are typically kept frozen during inference on real images, thus, they lead to poor generalization on real scenes which is called Sim-to-Real gap for IID tasks. In this work, we propose to solve both the challenges by using bank of Generative Adversarial Networks (GANs) as priors where each GAN is trained independently only on a single image component, thus alleviating the challenge of cross-contamination. Moreover, even though our GANs are trained on synthetic data, we allow for careful fine-tuning of the generator priors for inference on real scenes to bridge the Sim-to-Real gap.
Figure 1: Intrinsic Image Decomposition (IID) framework. We consider a commonly used image decomposition framework that aims at decomposing the natural image into its light independent (albedo), light dependent (shading), and optionally residual (specular) components. The forward model $f(\cdot)$ from the intrinsic components to the natural image is simply given by multiplication of albedo and shading with addition of specular as a residual (here, $sRGB(\cdot)$ indicates the standard tone-mapping operation). However, the inverse mapping of natural image to its intrinsic components is highly ill-posed inverse problem that we aim to solve.
The idea of using a GAN model as prior is explored widely for a variety of inverse imaging problems such as image inpainting (Yeh et al., 2016; Bora et al., 2017; Gu et al., 2019), HDR imaging (Niu et al., 2020), image super-resolution (Ledig et al., 2016; Bora et al., 2017; Gu et al., 2019), compressive sensing (Bora et al., 2017; Shah & Hegde, 2018a), and phase retrieval (Hand et al., 2018; Hyder et al., 2019), owing to their tremendous success in mapping a lower-dimensional latent distribution to the manifold of real images. Such approaches assume that the target image lies on the range of the GAN and attempt to obtain the corresponding latent code through optimization. However, unlike the imaging problems where the solution set consists of only a single natural image, in case of IID, the solution consists of several intrinsic components. For that reason, when used for IID task, such GAN prior is required to generate all the decomposition components from a single latent code. For example, in a previous work (Guo et al., 2020) all intrinsic components are generated simultaneously by a single GAN. Such a design is susceptible to the problems of poor disentanglement and cross-contamination as all components are generated by a single model and tied with a single latent code.
To this end, in this work we propose to use a bank of independently trained generators as a prior: we train three separate GANs – one each for albedo, shading, and specular. We model the natural image as a combination of the intrinsic components generated by these component-specific GANs, and find the latent codes for the albedo, shading, and specular GANs such that when input to the corresponding GAN model, they accurately generate their respective intrinsic component, and produce the input image when combined (Fig. 2). Since each GAN is trained independently on only a single image component, each generator has no capability of generating anything else other than the image component on which it is trained. Therefore, an independently trained GAN provides a much stronger constraint on the solution space as compared to the existing methods.
Using more than one GAN as a prior comes with new set of challenges. Existing approaches restrict the solution space by inverting the mapping of a pre-trained GAN, i.e. finding the latent code corresponding to the target image. Inverting the mapping of a GAN is a highly challenging and well-studied problem known as GAN Inversion. In our case, we need to obtain the latent codes for all the GANs in our bank – meaning inverting multiple GANs simultaneously. Since all the existing GAN inversion methods are limited to inverting only a single GAN, it is not straightforward to use them to jointly invert all the GANs in our bank. To this end, we propose a new Joint GAN Inversion algorithm, JoIN, that can jointly invert multiple GANs without requiring any component-wise supervision (see Fig. 2). To the best of our knowledge, JoIN is the first method to demonstrate successful inversion of multiple GANs at once.
Key to successful joint inversion is our kNN-based loss regularization that ensures the latent code estimates for each intrinsic component remains within the distribution manifold of its associated GAN. Further, we use only the synthetic data to train our GANs, and show that our approach successfully generalizes to real images by adapting generator fine-tuning technique from the existing GAN inversion literature in our pipeline: unlike the existing learning-based approaches that keep the priors frozen during the inference, we allow for a careful fine-tuning of our generator priors to bridge the Sim-to-Real gap. We evaluate our approach on different datasets such as faces, materials, and indoor scenes, and provide extensive comparisons on both the synthetic and real images.
Since we train a separate GAN model for each image component, our approach remains modular with respect to the choice of forward model, i.e. if we choose to add a new component (e.g. specular) to our existing forward model, we can reuse the GANs trained for existing components (albedo and shading) in our bank, and need to train only one GAN from scratch. We demonstrate such capability along with other ablation studies in our experiments.
Figure 2: Overview of our approach. Left: We use a bank of pre-trained GANs as a prior where each GAN is trained only on a single image component using synthetic data. We design the problem of decomposition as a joint GAN inversion problem on multiple GANs using the input image $I^{*}$ as the only supervision. Step 1: We aim to optimize the latent codes of each GAN in a way that after passing the outputs of individual GANs through forward mapping $f(\cdot)$, the resulting image estimate $\hat{I}$ resembles the input image $I^{*}$. Apart from using reconstruction loss, we propose to use kNN loss on the latent codes to strongly enforce the priors learnt by each GAN. Such approach leads to successful decomposition on synthetic data. Step 2: For decomposition on real images, we bridge the Sim-to-Real gap by carefully fine-tuning the individual GANs in a way that they can represent the real image features while still maintaining strong component-wise priors.
To summarize, our key contributions are following:
  • We propose to use a bank of independently trained GANs as priors to tackle the IID task. Training each GAN separately for individual intrinsic components provides much stronger priors and mitigates the challenges of poor disentanglement and signal cross-contamination.
  • We propose a new Joint GAN Inversion algorithm, JoIN, that can jointly invert multiple GANs without requiring any component-wise supervision (see Fig. 2(step 1)). Our method decomposes an image by successfully inverting it onto albedo GAN, shading GAN, and specular GAN simultaneously using only the natural image as a supervision. To the best of knowledge, ours is the first method to successfully invert multiple GANs at once.
@@
  • We introduce a novel kNN-based regularization term for joint GAN inversion allowing to better retain their distribution properties of each GAN as compared to existing regularizations.
  • We show that even though the bank of GANs is trained entirely on synthetic data, our approach can achieve excellent generalization on the real images by leveraging existing GAN inversion techniques such as generator fine-tuning (see Fig. 2(step 2)). Our model also offers flexibility of decomposing into either two components (albedo, shading) or three components (albedo, shading, specular).
@@ 2.1 Image Decomposition
The intrinsic decomposition of images is a long-standing problem that emerged with the retinex and lightness theory (Land & McCann, 1971; Horn, 1974). Using a purely diffuse model, early work computed shadow-free reflectance images of faces (Weiss, 2001) using time-lapse sequences. User-assisted methods were the first to successfully tackle the problem for single natural images (Tappen et al., 2003; Bousseau et al., 2009). With the rise of deep learning a large body of work emerged using neural networks to predict intrinsic images (Zhou et al., 2015; Nestmeyer & Gehler, 2017; Li & Snavely, 2018b; a; Yu & Smith, 2019; Li et al., 2020). While feed-forward methods such as (Das et al., 2022b; a; Yu & Smith, 2021) bring improvements in terms of model architecture and loss functions, they are prone to cross-contamination since they use a single model for predicting multiple components. Recently, CRefNet (Luo et al., 2023) proposed decoder-sharing transformer architecture for albedo prediction, but their model is limited to only a single component unlike our approach that can provide full decomposition of an image into 2 or 3 components. Similarly, S-Aware (Jin et al., 2023) too is limited to predicting only the albedo. Careaga & Aksoy (2023) predicts albedo and shading using two stage networks, but requires dense pseudo ground truth for generalization to real images. Liu et al. (2020) presented a method that shares some ideas with our pipeline: they build priors independently for shading and reflectance using image collections and adversarial auto-encoders. Contrary to our approach, all such feed-forward methods use a fixed forward model and would require full retraining if a component such as specular were to be added. One can refer to a recent review (Liu et al., 2024) for a summary of the research in intrinsic decomposition.
@@
For photo collections, many works explored this problem using geometry priors (Laffont et al., 2012; 2013; Duchêne et al., 2015; Philip et al., 2019; 2021), and recently methods were introduced working on videos (Ye et al., 2014; Bonneel et al., 2014; Meka et al., 2016; 2021). One of the main applications of IID lies in image relighting for which decomposing an image in specific components such as albedo, shading, shadows or normals is often an intermediate step both for faces (Pandey et al., 2021; Hou et al., 2021; Yeh et al., 2022) and outdoor picture (Griffiths et al., 2022). A recent work (Zhang et al., 2024) trains an end-to-end image relighting model by encoding the source scene and target illumination into a latent space, and recovers the albedo maps of the scene from its latent representation. However, the quality of intrinsic components obtained by such methods is generally poor since their primary aim is relighting, while intrinsic decomposition is treated as byproduct. In the context of the more constrained problem of material acquisition, methods often aim at recovering richer information. They usually predict a full SVBRDF map composed of albedo but also normals, roughness, and specular maps (Aittala et al., 2015; Deschaintre et al., 2018; 2019). Some recent methods proposed to use adversarial networks as part of their SVBRDF estimation pipelines (Zhou & Kalantari, 2021; Zhou & Khademi Kalantari, 2022). Closer in spirit to our method MaterialGAN (Guo et al., 2020) proposes to use a single GAN trained on all the channels of SVBRDF maps and to invert a latent code to recover an SVBRDF. Contrary to our method it requires several input images and known lighting conditions for each input.
Leveraging different learning-based priors for different signal components has been explored in previous works such as DoubleDIP (Gandelsman et al., 2019) and its extensions (Ren et al., 2020; Zhou et al., 2023; Zhuang et al., 2022) that use one untrained deep image prior (DIP) for each signal component. Their approach relies on strong internal self-similarity property, i.e., the patches within each component have higher self-similarity than the patches in the combined image, while the patches across the two components are highly dissimilar. This strategy has proven effective in tasks like image dehazing, layer separation, and watermark removal, where the components are often structurally dissimilar. However, the assumption of structural dissimilarity of the components may not hold for albedo and shading, since their features (shape, edges) and structure are highly similar to each other. Evidently, UIDNet (Zhang et al., 2022) demonstrated that applying DoubleDIP to IID results in severe cross-contamination and a loss of fine details. They introduce additional hand-crafted priors, and use encoder-decoder with skip connections to mitigate these issues. Using untrained DIP priors is computationally expensive with inference times of several minutes. In our case, instead of untrained DIPs, we employ multiple trained GANs that provide highly accurate priors, and by inverting them jointly, we effectively avoid the cross-contamination, preserve details, and achieve significantly faster inference times.
@@ 2.2 GAN Priors and GAN Inversion
Inverting a GAN to obtain the latent code corresponding to an image is a key step in solving inverse problems (Shah & Hegde, 2018b). Owing to its importance, many GAN inversion approaches have been recently proposed as reviewed in Xia et al. (2021). We can classify these methods in three broad classes: optimization-based, encoder-based, and hybrid.
@@
Optimization-based methods such as Lipton & Tripathi (2017b); Creswell & Bharath (2019); Zhu et al. (2016); Lipton & Tripathi (2017a); Karras et al. (2020); Abdal et al. (2019; 2020) regress a latent code using gradient descent by minimizing a loss between a target image and the generated image produced by the latent code. The key variations among these approaches are the latent space chosen for the optimization ($\mathcal{Z},\mathcal{W},$ and $\mathcal{W^{+}}$ space); the loss function(s) they use to calculate the similarity between the target and the estimate; and additional criterion used for aiding optimization.(such as regularization, improved initialization, and early stopping). Although these methods can produce relatively precise outcomes, their iterative optimization scheme is slower than forward methods and can converge to local minima when badly initialized.
@@
Contrary to optimization-based approaches, encoder-based methods propose to use an end-to-end framework based in which an encoder that takes the target image as input, predicts the latent code directly (Xia et al., 2021; Pidhorskyi et al., 2020; Richardson et al., 2021a; Alaluf et al., 2021; Chai et al., 2021a; Larsen et al., 2016; Zhu et al., 2016; Brock et al., 2016; Perarnau et al., 2016; Tov et al., 2021; Wei et al., 2021; Wang et al., 2021; Yao et al., 2022; Moon & Park, 2022; Mao et al., 2022). Encoder-based methods have faster inference times, but they suffer from lower result quality and poorer generalization. They can be used as initializer for optimization-based inversion. Methods that combine both optimization and encoder-based initialization are known as hybrid approach (Chai et al., 2021b; Bau et al., 2019c; Zhu et al., 2016; Alaluf et al., 2021; Bau et al., 2019b; a; Huh et al., 2020; Wei et al., 2021).
@@
Recently, a fourth category of approaches has emerged that advocates fine-tuning the generator weights after optimizing the latent code (Roich et al., 2021; Feng et al., 2022). Such methods produce better reconstruction since the generator is allowed to fit the target image but fine-tuning the generator can lead to distortions in the generator manifold losing important properties. More recently, (Bhattad et al., 2023) has shown that StyleGAN has an internal representation of intrinsic images that can be accessed by finding an appropriate offset to the latent codes. However, this method is not applicable for real images, since the offsets obtained are not compatible with the inverted latent codes of the real images.
In recent works, a new category of generative models known as diffusion models (Dhariwal & Nichol, 2021) are becoming increasingly popular in many application domains owing to their superior performance in generating high-quality and diverse images. Several works (Feng et al., 2023; Li et al., 2023) leverage diffusion models as priors in solving inverse problems such as reconstruction and super-resolution. Similar in spirit to our method, the concurrent work Wang et al. (2024) has explored using independently trained diffusion models for blind image deblurring, where a blurry image is decomposed into a clean image, blur kernel, and a tilt map. Note that unlike the IID, here the target components are structurally incoherent and dissimilar, e.g. the blur kernel is typically much lower dimensional and exhibit distinctly different features than the natural image. Moreover, Diffusion models are computationally more expensive than GANs during both training and inference, and does not allow for smooth transitions in latent space (that can be leveraged for image relighting in our case). While we acknowledge leveraging multiple diffusion priors for the IID task is a significant direction for future research, especially to model larger and more diverse datasets, our work demonstrates that GANs offer a powerful and efficient tool for IID, particularly when computational resources and inference speed are important considerations.
@@ 3 Our Approach
In this work, we propose to solve intrinsic image decomposition using an optimization formulation that leverages a bank of GANs as priors. We use the fact that each image component distribution can be closely approximated using a GAN model, and create a bank of GANs by training a separate GAN model for each image component that we desire to recover (See Fig. 2(left)). Since each GAN maps a lower dimensional latent space to an image component space, solving the inverse problem aims to recover the latent codes corresponding to each image component which can be formulated as the joint inversion of multiple GANs as shown in Fig. 2(step 1).
@@ 3.1 JoIN: Joint Inversion of Multiple GANs
For $G_{i}$ we follow the architecture of StyleGAN2 (Karras et al., 2020), while $w_{i}$ refers to an intermediate latent code in $\mathcal{W}$ space. Note that the StyleGAN-based architecture first maps a Gaussian distribution ($\mathcal{Z}$) to an intermediate latent space $\mathcal{W}$ using a non-linear mapping network, and $\mathcal{W}$-space is then mapped to distribution of real images. $\mathcal{W}$-space is known to provide higher degree of disentanglement and smoother control over various image features (Karras et al., 2019), thus has emerged as more suitable choice for performing GAN inversion (Xia et al., 2021). We follow the same and perform the joint inversion in $\mathcal{W}$-space. Our estimate of $I$ is thus given by:
$I(w_{1},\dots,w_{n})=f\left(T^{-1}_{1}\left(G_{1}\left(w_{1}\right)\right),\dots,T^{-1}_{n}\left(G_{n}\left(w_{n}\right)\right)\right)$ (3)
@@
Eq. 4 is analogous to the standard GAN inversion formulation, except for the fact that it aims to invert $n$ Generators simultaneously instead of one. The loss in Eq. 4 is minimized over all the latent codes $w_{i}$s simultaneously, and the final estimates $\widehat{w}_{i}$s can be passed to Eq. 2 to obtain the image components $C^{*}_{1\dots n}$. We use E-LPIPS loss (Kettunen et al., 2019) (robust version of LPIPS (Zhang et al., 2018) perceptual loss) as the loss function $\mathcal{L}$ unless otherwise specified.